jeannie has 10 books overdue at the library. she rembers she checked out 2 science books two weeks before she checked out 8 childeren's books. The daily fine per book is $0.20. If her total fine is $11.60, how long was each book over due?
Answer by ewatrrr(24785)


Hi,
Let d be the days the 8 books were over due. Then (d+14) will be for the science books
8*.20d + 2*.20(d+14) = $11.60
simplify and Solve
1.60d + .4d + 5.60 = $11.60
2d = 6.00
d = 3
3 days for the children's books
17 days for the science books
check your answer
8*.20*3 + .4*17 = 4.80 + 6.80 = $11.60